Write the expression for y as a function of x and
Answer #1

Amplitude A = 8.00 cm

wavelength \lambda = 80.0 cm

frequency f = 3.00 Hz

general equation for a sinusoidal wave travelling along the negative x axis

y = A \sin {(\frac{2\pi\,x}{\lambda}+2\pi\,f\,t+\phi)}

where \phi is the phase constant.

(a)

since y (0, t) = 0 for t = 0

\phi = 0

hence wave equation is y = (0.08 ) \sin {(\frac{2\pi\,x}{0.80}+2\pi\times\,3.00\,t)}

where A and \lambda are in meters.

(b)

since y (x, 0) = 0 for x = 10.0 cm

wave equation is y = (0.08 ) \sin {(\frac{2\pi\times0.10}{0.80}+2\pi\times\,3.00\times\,0+\phi)}

where A and \lambda are in meters.

so \phi=\frac{-\pi}{4}

hence wave equation is y = (0.08 ) \sin {(\frac{2\pi\,x}{0.80}+2\pi\times\,3.00\times\,t+\frac{-\pi}{4})}

where A and \lambda are in meters.